OpenStudy (anonymous):

The exact value of cos(-5pi/3) can be written in the form ((a sqr root b)/c) . Determine the values of a, b and c. I know that the angle is -300 therefore the reference angle is 60 in the first quadrant. Also, by looking at the special angles I know that b and c are 1/2, but will 'a' be positive or negative ?

OpenStudy (anonymous):

cos(-5pi/3) = -cos(5pi/3) = -cos(2 pi - pi/3) = -cos(pi /3) =- cos 60 =- 1/2 a=-1 b=1 c=2

OpenStudy (savvy):

dude cos(-x)=cos(x) hence \[\cos(-5\pi/3)=\cos(5\pi/3)\] \[= \cos(2\pi - \pi/3)\] \[=\cos(\pi/3)\] =1/2 hence, a and \[\sqrt{b}\] could take any values such that their product is 1. c=2

OpenStudy (anonymous):

Should be \[\cos (2\pi-5\pi/3)\]

OpenStudy (savvy):

where...????

OpenStudy (savvy):

no its in general...

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

sorry for that. I just though it in a wrong way